# MambaNUT: Nighttime UAV Tracking via Mamba and Adaptive Curriculum Learning

### v1

@@ 3.4 Vision Mamba for Tracking
where $E$ is the embedding dimension of each token. After obtaining the input tokens $t^{0}_{1:\mathcal{K}}$, we feed them into the encoding layer, where they are processed through stacked L layers of bidirectional Vision Mamba (Vim) encoders. Let $E^{l}$ denote the Vim layer at layer $l$, where forward propagation procedure involves all tokens from the layer $(l-1)$ via $t_{1:\mathcal{K}}^{l}=E^{l}(t_{1:\mathcal{K}}^{l-1})+t_{1:\mathcal{K}}^{l-1}$. The detailed structure of the bidirectional Vision Mamba encoders $E^{l}$ is illustrated on the right side of Fig. 3. The input $t^{0}_{1:\mathcal{K}}$ is first normalized and then processed separately through two distinct linear projection layers to obtain the intermediate features $\mathbf{V}$ and $\mathbf{Q}$:
@@
$Y$ $=Linear(\bm{V}_{forward}^{\prime})+Linear(\bm{V}_{backward}^{\prime}),$
@@
3.5 Tracking head and loss function
@@
$\mathcal{L}_{total}=\mathcal{L}_{cls}+\lambda_{iou}\mathcal{L}_{iou}+\lambda_{L_{1}}\mathcal{L}_{L_{1}}+\gamma\mathcal{L}_{ADB}$ (8)
@@
4 Experiment
In this section, we provide a thorough evaluation of our method using three nighttime UAV tracking benchmarks: NAT2024-1 [13], NAT2021 [67], and UAVDark135 [35]. Our evaluation is performed on a PC that was equipped with an i9-10850K processor (3.6GHz), 16GB of RAM, and an NVIDIA TitanX GPU. We evaluate our approach by comparing it with 16 state-of-the-art (SOTA) trackers, as detailed in Table 1.
Table 2: Comparison of precision (Prec.) and success rate (Succ.) between MambaNUT and the ten SOTA trackers on UAV123 [2].
Tracker MambaNUT(Ours) DCPT [73] AVTrack-DeiT [39] Sam-DA [14] Aba-ViTrack [38] MAT [71] TCTrack [6] UDAT-BAN [67] UDAT-CAR [67] SiamCAR [19] Ocean [70]
Prec. 86.5 85.2 84.8 76.1 86.2 86.7 77.3 76.1 75.5 76.0 78.3
Succ. 68.4 67.6 66.8 57.8 66.4 68.7 60.4 59.0 57.2 61.4 59.6
4.1 Implementation Details
Model Variants. We trained two variants of MambaNUT, each with different configurations, as described below:
  • MambaNUT-Tiny. Backbone: Vim-Tiny; Search region size: [256×256]; Template size: [128×128];
  • MambaNUT-Small. Backbone: Vim-Small; Search region size: [256×256]; Template size: [128×128];
Training. We use training splits from multiple datasets, including four daytime datasets: GOT-10k [25], LaSOT [10], COCO [43], and TrackingNet [47], and three nighttime datasets: BDD100K-Night, SHIFT-Night, and ExDark [46]. Notably, we select the images labeled as ”night” from the BDD100K [68] and SHIFT [52] datasets to construct the BDD100K-Night and SHIFT-Night. During training, the two variants of the tracker share the same training pipeline to maintain consistency and comparability. The batch size is consistently set to 32. We use the AdamW optimizer with a weight decay of $10^{-4}$, and an initial learning rate of $4\times 10^{-5}$. The total number of training epochs is fixed at 300, with 60,000 image pairs processed per epoch. The learning rate is reduced by a factor of 10 after 240 epochs.
@@ 4.1 Implementation Details
Figure 4: Qualitative evaluation on three video sequences from NAT2024-1: L05011, L07001, and L05015.
4.2 Overall Performance
@@
NAT2024-1: NAT2024-1 [13] is a long-term tracking benchmark featuring multiple challenging attributes, comprising 40 long-term image sequences with a total of over 70K frames. As shown in Table 1, our MambaNUT-Small tracker outperforms 16 state-of-the-art (SOTA) trackers on this benchmark, achieving a precision of 83.3%, a normalized precision of 76.9%, and a success rate of 63.6%. This surpasses the second-best tracker by 2.4%, 1.5%, and 1.5% in each metric, respectively. We also select three representative video sequences from NAT2024-1 for visualization in Fig. 4. As shown, MambaNUT-Small tracks the target objects more accurately than the seven SOTA trackers.
@@
NAT2021: NAT2021 [67] includes 180 testing videos, offering a challenging and large-scale benchmark for nighttime tracking. As shown in Table 1, MambaNUT-Small demonstrates competitive performance compared to the SOTA trackers. It achieves the highest precision and normalized precision, outperforming the previous top-performing tracker, DCPT, by more than 1.0% in both metrics, with only a slight 0.2% gap in success rate compared to DCPT.
@@
UAVDark135: The UAVDark135[35] benchmark consists of 135 test sequences and is widely used as the benchmark for nighttime tracking. From Table 1, MambaNUT-Small achieved a new state-of-the-art score of 70.0% in precision and 57.1% in success rate. Additionally, our MambaNUT-Tiny ranks third in all three metrics.
UAV123: UAV123 [2] is a large-scale aerial tracking benchmark with 123 challenging sequences and over 112K frames. To demonstrate that our proposed strategy of learning from easy (daytime) to hard (nighttime) significantly improves the generalization of the deep model, we compare our tracker with eight SOTA trackers on this daytime UAV tracking benchmark. Table 2 presents the Prec. and Succ. of the competing trackers on UAV123. MambaNUT ranks second, with only slight gaps of 0.2% in Prec. and 0.3% in Succ. compared to MAT.
4.3 Efficiency Comparison
@@
In Table 1, we also compare the inference speed on GPU, floating-point operations per second (FLOPs), and number of parameters of the proposed trackers with SOTA trackers to highlight the proposed superior efficiency. Notably, as AVTrack-DeiT feature adaptive architectures, the FLOPs and Params of it vary within a range, spanning from the minimum to the maximum values. As observed, although DCPT achieves comparable performance to our MambaNUT-Small, MambaNUT-Small runs in real-time at over 75 fps, i.e., more than twice DCPT’s speed, and uses only 1.1 GMacs and 4.1 million parameters, significantly less than DCPT’s 42 GMacs and 99 million parameters. While trackers like AVTrack-DeiT and Aba-ViTrack achieve higher tracking speeds than our method, their performance across multiple nighttime UAV tracking benchmarks is significantly lower. This comparison in terms of computational complexity also underscores the efficiency of our methods.
@@
Table 3: Illumination-Oriented Evaluation comparison with the 16 SOTA Trackers, evaluated on NAT2024-1[13].
LAI IV
Trackers Prec. Succ. Prec. Succ.
MambaNUT-Small(Ours) 0.877 0.672 0.772 0.556
DCPT[73] 0.854 0.658 0.693 0.513
AVTrack-DeiT[39] 0.787 0.594 0.662 0.463
TDA-Track[13] 0.773 0.519 0.504 0.328
Sam-DA[14] 0.852 0.571 0.600 0.404
SGDViT[62] 0.569 0.403 0.343 0.249
Aba-ViTrack[38] 0.824 0.628 0.627 0.464
HiT-Base[28] 0.668 0.512 0.470 0.346
MAT[71] 0.854 0.655 0.705 0.532
TCTrack++[7] 0.732 0.497 0.515 0.333
TCTrack[6] 0.786 0.514 0.593 0.358
UDAT-BAN[67] 0.750 0.543 0.500 0.323
UDAT-CAR[67] 0.722 0.524 0.568 0.395
HiFT[4] 0.642 0.455 0.390 0.262
SiamAPN++[5] 0.720 0.502 0.586 0.374
SiamCAR[19] 0.739 0.545 0.528 0.389
Ocean[70] 0.704 0.459 0.629 0.379
4.4 Illumination-Oriented Evaluation
@@
To further evaluate the performance of MambaNUT in nighttime scenarios, we conduct an analysis focused on the challenges of Low Ambient Illumination (LAI) and Illumination Variation (IV) on NAT2024-1. The evaluation results are shown in Table 3, and more attribute-based evaluation results are provided in the supplemental materials. As observed, our tracker significantly outperforms the SOTA trackers in these two attributes, achieving over a 2.0% improvement in both precision and success rate compared to the second-best tracker, with a remarkable 6.7% improvement in precision on the IV challenge.
4.5 Ablation Study
To validate the effectiveness of our proposed adaptive curriculum learning method, we conducted ablation studies using MambaNUT-Small as the baseline on the NAT2024-1 [13] dataset. Comprehensive results are provided in the supplementary material.
Figure 5: Visualization maps between without and with ACL.
@@ 4.5 Ablation Study
Table 4: Impact of Sampling Scheduler (SS) and Loss Scheduler (LS) on the performance of the baseline trackers on NAT2024-1.
Method SS LS Prec. Norm.Prec Succ.
MambaNUT-Small 79.6 73.8 60.6
$\checkmark$ 81.5↑1.9 75.3↑1.5 61.8↑1.2
$\checkmark$ $\checkmark$ 83.3↑3.7 76.9↑3.1 63.6↑3.0
@@
Impact of Adaptive Curriculum Learning (ACL) strategy: To validate the effectiveness of the proposed adaptive curriculum learning strategy, Table 4 presents the evaluation results on NAT2024-1, progressively incorporating two levels of curriculum schedulers, i.e., sampling scheduler (SS) and loss scheduler (LS), into the baseline. As observed, the incorporation of SS significantly enhances both Prec., Norm.Prec, and Succ. for the baseline tracker. With the further application of LS, the improvements become even more significant, with all increases exceeding 3.0%. Fig. 5 also demonstrates that by incorporating our ACL into the baseline tracker, more robust and discriminative feature representations are achieved, particularly enhancing the consistency of feature distribution across consecutive frames in long-term tracking. This comparison further demonstrates the effectiveness of our method in enhancing robust feature representations learning under low-light conditions using Mamba.
@@
Table 5: Performance comparison of different loss function scheduler selections.
@@
Impact of Loss Function Scheduler: To demonstrate the superiority of the proposed ADB loss in performance, we train separately MambaNUT-Small using Focal[42] and WCE [51] loss for comparison. The evaluation results on NAT2024-1 are shown in Table 5. From the table, while using Focal and WCE loss as the loss scheduler improves performance, the best precision improvement is only 2.2%, and the improvements in norm.precision and success rate remain below 2.0%, which is far behind our approach, where all three metrics show improvements above 3.0%.
Table 6: Impact of different sampling weights on the performance of baseline trackers on NAT2024-1.
Method $\theta$ Prec. Norm.Prec Succ.
MambaNUT-Small - 79.6 73.8 60.6
100 81.9↑2.3 75.1↑1.3 61.6↑1.0
150 83.3↑3.7 76.9↑3.1 63.6↑3.0
200 82.3↑2.7 75.6↑1.8 62.4↑1.8
Impact of Sampling Weight on Nighttime Data: In the proposed sampling scheduler, we set a constant $\theta$ (Seeing Eq (3)) to control the sampling weight of nighttime data as the training process progresses. We trained MambaNUT-Small using varied values of $\theta$ ranging from 100 to 200 with increments of 50. Evaluation results are presented in Table 6. As shown, our tracker achieves best performance when the constant is set to 150. These significant differences clearly highlight the substantial impact of nighttime data weight on tracking performance.
5 Conclusion
In this work, we propose MambaNUT, a novel Mamba-based nighttime UAV tracking framework that exploits Mamba’s exceptional ability to model long-range dependencies with linear complexity. Additionally, we incorporate an adaptive curriculum learning strategy into this framework by designing two curriculum schedulers for sampling and loss propagation. These schedulers dynamically guide the model to progress from imbalance to balance and from easy to hard across daytime and nighttime data. Extensive experiments demonstrate that our MambaNUT achieves state-of-the-art results on three nighttime UAV tracking benchmarks, while offering advantages in computational complexity and parameter efficiency.

### v2

@@ III-D Vision Mamba for Tracking
where $E$ is the embedding dimension of each token. After obtaining the input tokens $t^{0}_{1:\mathcal{K}}$, we feed them into the encoding layer, where they are processed through stacked L layers of bidirectional Vision Mamba (Vim) encoders. Let $E^{l}$ denote the Vim layer at layer $l$, where forward propagation procedure involves all tokens from the layer $(l-1)$ via $t_{1:\mathcal{K}}^{l}=E^{l}(t_{1:\mathcal{K}}^{l-1})+t_{1:\mathcal{K}}^{l-1}$. The detailed structure of the bidirectional Vision Mamba encoders $E^{l}$ is illustrated on the right side of Fig. 3. The input $t^{0}_{1:\mathcal{K}}$ is first normalized and then processed separately through two distinct linear projection layers to obtain the intermediate features $\bm{V}$ and $\bm{Q}$:
@@
$\bm{Y}$ $=Linear(\bm{V}_{forward}^{\prime})+Linear(\bm{V}_{backward}^{\prime}),$
@@
III-E Tracking head and loss function
@@
$\mathcal{L}_{total}=\mathcal{L}_{cls}+\lambda_{iou}\mathcal{L}_{iou}+\lambda_{L_{1}}\mathcal{L}_{L_{1}}+\gamma\mathcal{L}_{ADW}$ (8)
@@
IV Experiment
In this section, we provide a thorough evaluation of our method using three nighttime UAV tracking benchmarks: NAT2021 [13], NAT2024-1 [14], and UAVDark135 [45]. Our evaluation is performed on a PC that was equipped with an i9-10850K processor (3.6GHz), 16GB of RAM, and an NVIDIA TitanX GPU. We evaluate our approach by comparing it with 16 state-of-the-art (SOTA) trackers, as detailed in Table I.
IV-A Implementation Details
Model. We use Vision Mamba, i.e., Mamba®-Small [57], as the backbone to build our MambaNUT tracker for evaluation. The head of MambaNUT consists of a stack of four Conv-BN-Relu layers, with the search region and template sizes set to 256 × 256 and 128 × 128, respectively.
Training. We use training splits from multiple datasets, including four daytime datasets: GOT-10k [8], LaSOT [7], COCO [58], and TrackingNet [9], and three nighttime datasets: BDD100K-Night, SHIFT-Night, and ExDark [25]. Notably, we select the images labeled as ”night” from the BDD100K [26] and SHIFT [24] datasets to construct the BDD100K-Night and SHIFT-Night. The batch size is consistently set to 32. We use the AdamW optimizer with a weight decay of $10^{-4}$, and an initial learning rate of $4\times 10^{-5}$. The total number of training epochs is fixed at 300, with 60,000 image pairs processed per epoch. The learning rate is reduced by a factor of 10 after 240 epochs.
@@ IV-A Implementation Details
Figure 4: Qualitative evaluation on two video sequences from NAT2024-1: L50001 and L50011.
IV-B Overall Performance
@@
NAT2024-1: NAT2024-1 [14] is a long-term tracking benchmark featuring multiple challenging attributes, comprising 40 long-term image sequences with a total of over 70K frames. As shown in Table I, our MambaNUT tracker outperforms 16 state-of-the-art trackers on this benchmark, achieving a precision of 83.3%, a normalized precision of 76.9%, and a success rate of 63.6%, resulting in improvements of 2.4%, 1.5%, and 1.5% over the second-best tracker on these three metrics, respectively. We also select two representative video sequences from NAT2024-1 for visualization in Fig. 4. As shown, MambaNUT tracks the target objects more accurately than the 4 SOTA trackers.
@@
NAT2021: NAT2021 [13] includes 180 testing videos, offering a challenging and large-scale benchmark for nighttime tracking. As shown in Table I, MambaNUT demonstrates competitive performance compared to the SOTA trackers. It achieves the highest precision (70.1%) and normalized precision (64.6%), outperforming the previous top-performing tracker (i.e., DCPT) by more than 1.0% in both metrics, with only a slight 0.2% gap in success rate.
@@
UAVDark135: UAVDark135[45] benchmark consists of 135 test sequences and is widely used as the benchmark for nighttime tracking. From Table I, MambaNUT achieved a new SOTA score of 70.0% in precision and 57.1% in success rate, with a slight gap in success rate compared to DCPT.
IV-C Efficiency Comparison
@@
In Table I, we also compare our MambaNUT with SOTA trackers in terms of GPU inference speed, floating-point operations per second (FLOPs), and the number of parameters (Params.) to highlight its superior trade-off between performance and efficiency. Notably, as AVTrack-DeiT features adaptive architectures, its FLOPs and Params. vary within a range, spanning from the minimum to the maximum values. As observed, although DCPT achieves comparable performance to our MambaNUT, our method runs in real-time at over 75 fps, more than twice the speed of DCPT, and requires only 1.1 GMac FLOPs and 4.1 million parameters, significantly lower than DCPT’s 42 GMacs and 99 million. While trackers like AVTrack-DeiT and Aba-ViTrack achieve higher tracking speeds than our method, their performance across multiple nighttime UAV tracking benchmarks is significantly inferior. This comparison in terms of computational complexity also underscores the efficiency of our methods.
@@
Figure 5: Illumination-oriented evaluation comparison with the 8 SOTA trackers, evaluated on NAT2024-1[14].
IV-D Illumination-Oriented Evaluation
@@
To further evaluate the performance of MambaNUT in nighttime scenarios, we conduct an analysis focused on the challenges of Low Ambient Illumination (LAI) and Illumination Variation (IV) on NAT2024-1. Note that we also evaluated MambaNUT without the proposed ACL strategy, referred to as MambaNUT* for comparison. The precision plots are shown in Fig. 5. As observed, our tracker significantly outperforms SOTA trackers in these two attributes, achieving impressive 2.3% and 6.6% improvements in precision on LAI and IV, respectively, compared to the second-best tracker. Notably, incorporating the proposed ACL method leads to substantial improvements of 5.2% and 10.6% in precision for LAI and IV, respectively, over MambaNUT*, underscoring its effectiveness.
IV-E Ablation Study
Figure 6: Visualization of feature maps generated by MambaNUT without (middle) and with (bottom) the proposed ACL, with the first row displaying input search images from two sequences.
@@ IV-E Ablation Study
TABLE II: Impact of Fine-Tuning (FT), Sampling Scheduler (SS), and Loss Scheduler (LS) on baseline tracker performance on NAT2024-1.
Method Prec. Norm.Prec Succ.
baseline 79.6 73.8 60.6
baseline+FT 80.3↑0.7 74.6↑0.8 61.2↑0.6
baseline+FT+LS 81.3↑1.7 75.1↑1.3 61.9↑1.3
baseline+SS 81.5↑1.9 75.3↑1.5 61.8↑1.2
baseline+SS+LS 83.3↑3.7 76.9↑3.1 63.6↑3.0
@@
Impact of Adaptive Curriculum Learning (ACL) strategy: To validate the effectiveness of the proposed adaptive curriculum learning strategy, Table II presents the evaluation results on NAT2024-1, progressively incorporating two levels of curriculum schedulers, i.e., sampling scheduler (SS) and loss scheduler (LS), into the baseline. Notably, we make additional attempts by first training the model on daytime data and then fine-tuning (FT) it on nighttime data for 50 epochs. As observed, while FT on the daytime foundation tracker enhances performance, it only achieves results comparable to those with our SS. With the additional application of LS, the improvements become even more significant, with all three metrics gains exceeding 3.0%. Fig. 6 also demonstrates that by incorporating our ACL into the baseline tracker, more robust and discriminative feature representations are achieved, particularly enhancing the consistency of feature distribution across consecutive frames in long-term tracking. This comparison further demonstrates the effectiveness of our method in enhancing robust feature representations learning under low-light conditions using Mamba.
@@
TABLE III: Impact of different loss function schedulers on the performance of MambaNUT.
@@
Impact of Loss Function Scheduler: To demonstrate the superiority of the proposed ADW loss in performance, we train separately MambaNUT using Focal[59] and WCE [60] loss for comparison. The evaluation results on NAT2024-1 are shown in Table III. From table, while using Focal and WCE loss as the loss scheduler improves performance, the best precision improvement is only 2.2%, and the improvements in norm.precision and success rate remain below 2.0%, which is far behind our approach, where all three metrics show improvements above 3.0%.
Figure 7: Real-world test on an embedded device. Frame-wise performance is illustrated using CLE plots, with errors below the green dashed line (CLE = 20 pixels) considered successful tracking results.
V Real-world Tests
As shown in Fig. 7, we conducted real-world testing by deploying MambaNUT on a standard UAV platform equipped with a NVIDIA Jetson Orin NX to validate its performance. As shown, the main challenges include partial occlusion, camera motion, and background cluster. Nevertheless, MambaNUT exhibits robust performance, with all test frames maintaining a CLE below 20 pixels. Additionally, MambaNUT runs in real-time over 30 frames per second. Real-world testing results demonstrate that MambaNUT is well-suited for edge deployment on UAV platforms, achieving robust tracking performance in complex nighttime circumstances.
VI Conclusion
In this work, we propose MambaNUT, a novel Mamba-based nighttime UAV tracking framework that exploits Mamba’s exceptional ability to model long-range dependencies with linear complexity. Additionally, we integrate an adaptive curriculum learning (ACL) strategy into the framework with two schedulers for sampling and loss backward propagation. The proposed scheduler guides the model from imbalance to balance and from easy to hard across daytime and nighttime data. The Adaptive Data Weighted (ADW) loss scheduler employs a weighting scheme based on the size of training data and the IoU of individual instances. Extensive experiments demonstrate that our MambaNUT achieves state-of-the-art results on three nighttime UAV tracking benchmarks, while offering advantages in computational complexity.